OpenStudy (anonymous):

Solve the system of linear equations. 2x - 8y = 0 3x - 8y = 12

OpenStudy (anonymous):

2x - 8y = 0 2x = 8y x = 4y substitute to 3x - 8y = 12 3x - 8y = 12 3(4y) - 8y = 12 12y - 8y = 12 4y = 12 y = 12/4 y =3

OpenStudy (anonymous):

get it @kk54 ?

OpenStudy (anonymous):

So if its asking for the ordered pair example, 3,2 what would it be

OpenStudy (anonymous):

you mean x = 3 and y = 2 ??

OpenStudy (anonymous):

I'm not sure. My answer box just says enter an ordered pair like (3,2)

OpenStudy (anonymous):

hold on we got y = 3, just substitute it to 2x - 8y = 0 to get x 2x - 8y = 0 2x - 8(3) = 0 2x - 24 = 0 2x = 24 x = 24/12 x = 2

OpenStudy (anonymous):

x = 2, and y = 3 or (2,3)

OpenStudy (anonymous):

get it @kk54 ??

OpenStudy (anonymous):

thank you